### v1

# SpotLight: Shadow-Guided Object Relighting via Diffusion

Abstract
Recent work has shown that diffusion models can be used as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. Unlike typical physics-based renderers, however, neural rendering engines are limited by the lack of manual control over the lighting setup, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise lighting control can be achieved for object relighting simply by specifying the desired shadows of the object. Rather surprisingly, we show that injecting only the shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training. Specifically, we demonstrate its use with two neural renderers from the recent literature. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting.
| Unconditioned | Shadow-conditioned (top: shadow mapping, bottom: hand-drawn scribbles) composites from SpotLight | |||||
1 Introduction
Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process. When one is given a single photograph, however, no such control is possible. Relighting then becomes a hard task because of the complex surface interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture. The recent rise of powerful image generation models (e.g., [52]) has enabled exciting possibilities for image relighting [32, 73].
This paper focuses on providing local lighting control, where one wishes to modify the appearance of a single object. This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image. In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
In this paper, we present SpotLight, a method for realistic, controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than training a model specifically for the task of object relighting, our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow to a general-purpose neural renderer. Indeed, we find that simply specifying the desired lighting through the target shadow makes the renderer relight the object accordingly! SpotLight: Shadow-Guided Object Relighting via Diffusion shows two examples of such shadow-based lighting control: 1) the shadow of the couch is rendered using basic shadow mapping [19] according to the desired direction (top); and 2) one can even draw scribbles by hand to depict the desired shadow of the ottoman (bottom). To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained neural renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
The core technical contribution lies in our shadow injection approach which is robust to artifacts in the target shadow map. We build upon existing object compositing networks [75, 72] that operate on intrinsic images and exploit diffusion models’ learned priors to render realistic outputs. We demonstrate SpotLight handles complex scenarios and maintains physical plausibility without requiring extensive manual intervention nor any additional training.
In summary, our contributions are as follows. First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring fine-tuning or additional training. We demonstrate superior quality, both perceptually through a user study and quantitatively, compared to competitive specialized models. Finally, we release a novel evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
2 Related work
Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [58, 28, 59, 52]. These models offer control mechanisms for various modalities, such as text [48, 52], image content [53, 21], intrinsic images [75, 42, 72], classifier guidance [16] and cross-modal data [12, 69]. Notably, training-free guidance methods [43, 20, 3, 4] have enabled image editing capabilities without specific training, forming the basis of our approach.
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for style transfer [25, 41] or harmonization [51, 46, 35, 61] do not prioritize physically accurate shading of the objects. Recently, [47] utilized a single-view multi-illumination dataset [44] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [65, 11] for relighting indoor scenes. IC-Light [73] enforces consistency in appearance to relight portraits and various objects. Some methods specialize in outdoor scenes, using geometric priors like depth [27] or normalized coordinates [32] or both [38] to improve global lighting estimation. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13]. ZeroComp [75] learns to composite virtual objects in a zero-shot fashion, but does not offer lighting control. RGBX [72] uses diffusion models for material- and lighting-aware composition with text-guided control. Neural Gaffer [30] and DiLightNet [71] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. We bridge this gap by introducing a shadow guidance for object relighting.
Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 37, 76], though they generally lack controllability. Parametric models offer more control [23, 63, 15], but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
Shadow generation Generative models for rendering realistic shadows [74, 39, 29, 40], or harmonizing rendered shadows with the image [62] have been proposed. Other methods enable controllable shadows for 3D [55] and 2D objects [56, 57]. Our approach can leverage existing shadow generation methods and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
Image intrinsics Decomposing images into albedo and shading has long been studied [7, 34, 45, 64, 13]. Many works have also focused on estimating scene geometry, for example, depth [49, 8, 70, 31, 66, 67] and normals [18, 6, 50, 5, 68]. Recently, approaches such as [37, 76] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 31, 68, 2, 42]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
3 SpotLight
Diffusion-based neural renderers are currently limited by their inability to directly control shading, especially for specific objects as they are generally conditioned on albedo, normals, depth, and global shading. SpotLight addresses this limitation by enabling control over a single object lighting effects by defining a coarse shadow. Leveraging priors from pre-trained models and the diffusion framework, we introduce a method capable of harmonizing the object realistically from a provided shadow region. Fig. 2 presents an overview of the method, with each component described in detail below.
3.1 Rendering from intrinsics
Our method leverages a pre-trained diffusion-based neural renderer (e.g., ZeroComp [75], RGBX [72]) which accepts a multi-channel input of intrinsic maps —typically: albedo, shading, surface normals and depth, although other maps such as roughness and metallic can also be used—and outputs an RGB image consistent with these inputs. Our approach also requires shadow and object masks, and resp., where 1 indicates if the pixel is in shadow/object, and 0 otherwise. The discussion below assumes is based on latent diffusion [52], which performs the diffusion process in a low resolution latent space defined by a VAE encoder .
3.2 Blending shadows with the background
Provided a user defined shadow mask (), we first need to ensure that the diffusion model will properly blend the shadow in its final render. To achieve a perceptually pleasing shadow, we take inspiration from Blended Latent Diffusion [4] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadows composited over the image, . In practice, we obtain by compositing the object albedo and the shadow on the background. The update made at each timestep to incorporate the desired shadow latents is therefore
where is the shadow latent weight and is the shadow mask, first dilated by a square kernel to maintain the desired shadow softness on the edges, then bilinearly downsampled to the latent resolution (). Noise is added to the encoded composite using the DDIM [59] noise scheduler. Different from [4], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
3.3 Enhancing lighting control
Simply running a neural renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec. 4.6). We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by . Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample. We obtain the outputs of the diffusion renderer, which are typically -predictions [54], that is, a linear combination of the noise and the estimated image. We then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After applying eq. 2, we run a regular DDIM sampling step to obtain the latents .
3.4 Diffusion-based neural renderer backbone
Since SpotLight only applies changes to the latent space of a diffusion model, it can readily be applied to pre-trained diffusion-based neural renderers without any finetuning.
ZeroComp [75] provides a maskable shading map as input to a ControlNet-based neural renderer. Here, we use their approach directly and mask out the shading on both the object and the shadow region.
RGBX [72] provides a checkpoint of their XRGB model finetuned for inpainting rectangular masks. We mask out the bounding box including the object and shadow region.
At each denoising step, the denoising network is ran to obtain
where represents the network -prediction, represents the composited intrinsics. For ZeroComp, those intrinsics correspond to albedo, normals, depth, and partial shading and for RGBX, we use the albedo, normals, metallic, roughness and masked image.
3.5 Obtaining the final render
After having performed the denosing steps, we obtain an image by decoding the latents into by using the VAE decoder. The final result is obtained with the background preservation strategy proposed in [75], which obtains a shadow matte by computing the ratio of the image obtained with and without shadow guidance. This ensures the diffusion process does not affect the background image.
4 Evaluation on 3D object compositing
In this section, we quantitatively evaluate SpotLight against other methods that support lighting control on composition of 3D objects in 2D images. Additionally, we conduct a user study to assess human perception, comparing its performance with these competing approaches.
4.1 Evaluation dataset
We start from the evaluation dataset proposed in [75], where high-quality 3D models from the Amazon Berkeley Objects dataset [14] are inserted into background images cropped with a 50° field of view from the Laval Indoor HDR dataset [22]. The authors of [75] graciously provided us with reproducible Blender scene information, which we extend in two ways. First, we improve the shadow catcher using back-projected background depth estimated using Depth Anything V2 [67]. Second, object intrinsics are rendered with anti-aliased alpha mattes, allowing for more realistic compositing results. We use a depth-warping strategy on the HDR panorama as in [75] to account for spatially varying lighting around the object position. The dataset is rendered using the physically-based Cycles renderer with 2048 samples per pixel, a noise threshold of 0.01, and denoised using OpenImageDenoise [1]. We produce two versions of the dataset, each intended for different purposes.
“Reference-based” dataset This version identifies the dominant light source in the scene, extracted from the ground-truth warped panorama using the light extraction algorithm from [23]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3. We filter out samples where the background depth estimation failed. This data set is intended for quantitative evaluation (sec. 4.4), allowing the simulated ground truth of the inserted object to be used to compute metrics.
“User-controlled” dataset This second dataset simulates a scenario where a user specifies a desired light position. In this setup, five lighting directions are defined from left to right in increments in azimuth, with the light source passing behind the object to ensure the shadow remains visible. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,050 images. This dataset will be used in the user study (sec. 4.5) to assess the perceptual accuracy of renders in edited lighting scenarios.
| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) | Simulated GT |
4.2 SpotLight specifics
As mentioned in sec. 3.4, our proposed SpotLight readily applies to existing diffusion-based neural renderers. Here, we experiment with two approaches from the recent literature: ZeroComp [75] and RGBX [72].
ZeroComp backbone We train ZeroComp following the procedure outlined in [75], and use the same conditioned intrinsic estimators at inference time: StableNormal [68] for normals, IID [33] for albedo, and ZoeDepth [8] for depth. The masked shading in the background is obtained by dividing the image by the albedo and applying the same masking strategy. Given its superior results, we adopt this method as our default model for 3D object compositing.
RGBX backbone We use the RGBX model to extract intrinsic properties (normals, metallic, roughness, and albedo). We observe that results are generally overly bright, likely due to a domain gap between the training and evaluation images. To address this, we re-expose the background by a factor of 2 before feeding it to the network, then divide the output by 2, in linear space. We apply the same background preservation strategy as in ZeroComp.
Light control SpotLight requires a coarse shadow for guidance. Therefore, we use a real-time rasterization renderer (Blender’s EEVEE) to compute a simple shadow of the 3D object. This shadow is generated from a simple parametric distant sphere light and projected onto the mesh constructed from the estimated depth map of the background.
4.3 Baselines
We compare SpotLight against recent diffusion-based methods that support light control: DiLightNet [71], IC-Light [73], and Neural Gaffer [30], using their public implementations. While these baselines can relight objects, none are capable of generating a shadow around the object. For fairness, the physically-based renderer Cycles is therefore used to cast a realistic shadow. OpenImageDenoise [1] is also used to denoise the results of all methods. We begin by describing how light control is given to baselines, then proceed to describe per-method specifics.
4.3.1 Light control
In order to ensure fairness in evaluation, we aim to provide each method with a light parametrization that is as close as possible to the ground truth dominant light direction available in the evaluation dataset.
Both DiLightNet [71] and Neural Gaffer [30] require an environment map for lighting. We construct such an environment map with radiance along direction defined as the sum of a spherical gaussian and a constant term
where and are the RGB colors of the light and ambient terms resp., is the dominant light source direction, and is the bandwidth. is obtained by computing the average color over the background image. Panoramas from the Laval Indoor HDR dataset (excluding those in our test set) are used to estimate a single average intensity of the dominant light source, , which is set to the ratio of the integral of the brightest pixels in the panoramas, divided by the integral over all pixels. We further divide this integral by 2 to consider a single hemisphere and avoid overly bright dominant light sources. The light color is defined by , where is the normalized ambient color . We found that the bandwidth parameter did not advantage any specific method and therefore fixed it to , which we typically observe for an indoor light.
4.3.2 Per-method details
DiLightNet [71] DiLightNet requires a prompt describing the object, which we obtain by feeding the ground truth object composited on a black background to BLIP-2 [36] with the prompt “Putting aside the black background, this object is a”. To generate the results, depth and radiance hints are required. We supply DiLightNet with depth information from the 3D model of the object and render the radiance hints using the input environment map (sec. 4.3.1) as outlined in [71]. The rendered object is then composited with the background, along with the rendered shadow.
Neural Gaffer [30] We provide the method with the environment map described in sec. 4.3.1. The ground truth object is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at lower resolution (), its result is upsampled to then composited on the full resolution background. An erosion using a kernel is applied to the ground truth mask to mitigate boundary artifacts.
IC-Light [73] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning. We provide a rendering of the object lit by the ground truth environment map, giving a significant advantage to the method. No prompts are used.
4.4 Experimental results
Tab. 1 presents quantitative results comparing the various baselines (c.f. sec. 4.3) with two versions of SpotLight (c.f. sec. 4.2) on our “reference-based” dataset, where simulated ground truth is available (sec. 4.1). Here, all methods are conditioned on the dominant light direction using their specific light parametrization (sec. 4.3.1). It is worth noting that all three baselines (DiLightNet, IC-Light and Neural Gaffer) are given as input the ground truth object foreground and shadowed background, giving them a significant advantage. Fig. 3 shows qualitative results for each method. We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| DiLightNet [71]∗ | 25.36 | 0.950 | 0.059 | 0.024 | 0.044 |
| IC-Light [73]∗ | 28.59 | 0.965 | 0.043 | 0.018 | 0.034 |
| Neural Gaffer [30]∗ | 29.64 | 0.963 | 0.037 | 0.016 | 0.038 |
| SpotLight (RGBX) | 26.44 | 0.956 | 0.052 | 0.018 | 0.041 |
| SpotLight (ZeroComp) | 30.68 | 0.973 | 0.033 | 0.012 | 0.031 |
4.5 User study
A user study was conducted to evaluate the realism of local lighting control. Here, we use the “user-controlled” version of the dataset (see sec. 4.1) to evaluate whether the realism of results obtained by conditioning on a specific lighting direction, even if it is not the dominant light direction.
We evaluate the three methods with the best PSNR from tab. 1, namely SpotLight (with the ZeroComp backbone), Neural Gaffer [30] and IC-Light [73]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 40 random images are chosen, resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, as well as the left-right order. Examples are shown in fig. 5.
The user study was completed by observers. Following Giroux et al. [26], we used the Thurstone Case V Law of Comparative Judgement [60] to obtain a scaling -score for each method, where a higher scale indicates a higher human preference. The results, shown in fig. 4 show that our method is the favorite. The 95% confidence interval error bars also show that the margin in performance is statistically significant, indicating that SpotLight achieves the highest realism.
| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) |
4.6 Effect of parameter selection
Our method employs parameters that control the degree of shadow guidance ( in eq. 2) and the shadow blending ( in eq. 1). Tab. 2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the simulated ground truth is available (c.f. sec. 4.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 6 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible (fig. 7).
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| , (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.031 |
| (no guidance) | 31.65 | 0.976 | 0.030 | 0.011 | 0.029 |
| 28.60 | 0.966 | 0.044 | 0.015 | 0.036 | |
| 30.78 | 0.974 | 0.033 | 0.012 | 0.030 | |
| 29.32 | 0.970 | 0.039 | 0.014 | 0.034 |
| No guidance () | (ours) |
| (no blending) | (ours) |
5 Extensions
This section highlights additional capabilities derived from our framework to further demonstrate its versatility.
Shadow control on 2D object composition We demonstrate that SpotLight can be applied to 2D images by leveraging a ZeroComp [75] backbone trained without depth maps, using the normals [68] and albedo [33] estimators. Without access to a 3D model, we cannot rely on a rendering engine to cast shadows. Instead, we employ PixHt-Lab [57] to generate realistic soft shadows from a controllable point light coordinate. Here, a sample without shadow is provided for negative guidance , see sec. 3.3. To estimate the required pixel height of the object cutout, we attribute each pixel -coordinate as the pixel height for demonstration purposes, more sophisticated methods such as [56] could be used. After rendering with the backbone, we enhance the target by adding the high-frequency details from the source image. This is achieved by subtracting a Gaussian-blurred version of the source image ( = 1) from itself (as in unsharp masking) and adding the resulting details to the target. Example 2D compositing results are shown in fig. 8, which demonstrate realistic 2D object relighting from a user-controlled shadow.
| Object | Background | Light-controllable composites | ||
Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and pixel height in 2D. However, one can simply draw the desired shadow! Fig. 9 and SpotLight: Shadow-Guided Object Relighting via Diffusion (first row) show that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
6 Discussion
We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without requiring any additional training. Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
While the model can realistically estimate relighting from a shadow, generating a shadow map from scratch remains challenging. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
Acknowledgements
This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all members of the lab for discussions and proofreading help.