### v2

# SpotLight: Shadow-Guided Object Relighting via Diffusion

Abstract
Recent work has shown that diffusion models can serve as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise lighting control can be achieved for object relighting simply by providing a coarse shadow of the object. Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting. We also demonstrate other applications, such as hand-scribbling shadows and full-image relighting, demonstrating its versatility.
Background image | Previous method | SpotLight enables shadow-guided lighting control |
1 Introduction
Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process. When one is given a photograph, however, no such control is possible. Relighting then becomes a hard task because of the complex interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture. The recent rise of powerful image generation models (e.g., [54]) has enabled exciting possibilities for image relighting [31, 76].
This paper focuses on providing lighting control to recent diffusion image relighting frameworks, where one wishes to modify the appearance of a single object. This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image. In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
We present SpotLight, a method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [74, 29]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer. Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it to relight the object accordingly. We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object as it moves around it.
Fig. 1 shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [19] according to the desired lighting direction. From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting. To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
Our core technical contribution lies in our shadow guidance approach, which refines the guiding shadow input while realistically relighting the virtual object. Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control. This builds upon existing object compositing networks [78, 75] that operate on intrinsic images along with diffusion models’ learned priors to render realistic outputs. We demonstrate that SpotLight’s guidance mechanism handles complex scenarios and maintains physical plausibility without requiring manual intervention nor any additional training.
In summary, our contributions are as follows. First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring additional training. We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models. Finally, we release an evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
2 Related work
Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [59, 27, 60, 54]. These models offer control mechanisms for various modalities, such as text [50, 54], image content [55, 21], intrinsic images [78, 43, 75, 39], classifier guidance [16] and cross-modal data [12, 72]. Notably, training-free guidance methods [44, 20, 3, 4, 26] have enabled image editing capabilities without specific training, forming the basis of our approach.
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for harmonization [53, 48, 34, 62] do not prioritize physically accurate shading of the objects. Recently, [49] utilized a single-view multi-illumination dataset [45] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [67, 11] for relighting indoor scenes. IC-Light [76] enforces consistency in appearance to relight portraits and various objects. LumiNet [68] and ScribbleLight [14] achieve image relighting through illumination transfer or user-defined scribbles, respectively. Some methods specialize in outdoor scenes, using geometric priors like depth [25] or normalized coordinates [31] or both [40]. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13]. Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [78, 75, 39], which can be leveraged to achieve zero-shot object composition. However, these methods have limited lighting control. ZeroComp [78] provides no lighting control, RGBX [75] can only provide lighting control through a text prompt, and DiffusionRenderer [39] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set. Other diffusion-based approaches, such as Neural Gaffer [29], DiLightNet [74] and IllumiNeRF [79] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 36, 80, 47, 38], though they generally lack controllability. Parametric models [23, 65, 15] or GAN inversion methods [64] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
Shadow generation Generative models for rendering realistic shadows [77, 41, 28, 42], or harmonizing rendered shadows with the image [63] have been proposed. Other methods enable controllable shadows for 3D [56] and 2D objects [57, 58]. Our approach can leverage existing shadow generation methods (e.g., [58], see supp.) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
Image intrinsics Decomposing images into albedo and shading has long been studied [7, 33, 46, 66, 13]. Numerous works have also focused on estimating scene geometry, for example, depth [51, 8, 73, 30, 69, 70] and normals [18, 6, 52, 5, 71]. Recently, approaches such as [36, 80] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 30, 71, 2, 43]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
3 Background: diffusion renderers
Recent work [75, 78] has shown that diffusion models can be adapted to generate images from intrinsic maps including materials (e.g., albedo, roughness, metallic), geometry (e.g., surface normals and depth), and shading. This capability was built using a ControlNet with ZeroComp [78] or full finetuning with RGBX [75]. We refer to this new class of conditional generation models as “diffusion renderers”.
In these models, the denoising network is ran at each denoising step to obtain
where represents the network -prediction, the noisy latents, and the conditioning intrinsic maps.
It has been shown that these methods can realistically composite virtual objects into images in a zero-shot fashion, despite not being trained for this task. Provided a mask of the virtual object and its intrinsic maps (rendered with graphics shaders), the object can be trivially composited in most intrinsic maps of the background image (estimated with trained estimators). Since computing the virtual object shading requires complex lighting simulations, diffusion approaches estimate the shading by providing a partial shading map [78], or train inpainting models [75]. However, unlike traditional rendering, these methods offer no control over the lighting conditions.
In this work, we use a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, analogous to a spotlight moving around the object. We now describe our approach in detail.
4 SpotLight
We employ a user-provided, guiding shadow to steer the pre-trained diffusion process, thereby enforcing matching illumination on the inserted object. Our key insight is that even an approximate shadow encapsulates important lighting information. By integrating this cue early in the diffusion process, we simultaneously refine both the composite appearance and its shadow, yielding a natural blend with the scene.
4.1 Approach overview
Our approach requires two inputs: an object mask and a user-specified guiding shadow . This shadow, computed via fast rasterization or specified through hand-drawn scribbles, is used to guide the latent diffusion process. Without retraining the pre-trained renderer , we steer its generative process by incorporating the guiding shadow into the latent space.
4.2 Blending shadows
To fuse the guiding shadow with the latent representation, we update the noisy latent code at each diffusion timestep, similar to [4]. In our framework, the updated latent
serves as a soft constraint that steers the process toward a shadow-consistent output. Here, is the composite of the object albedo, the target shadow, and the background, and controls the strength of the guiding shadow. The latent shadow mask is obtained by downsampling the target shadow to the latent resolution (see supp. for details).
4.3 Shading the object via dual-branch guidance
A key element of our approach is to amplify the impact of the guiding shadow on the object shading through classifier-free guidance (CFG) [26]. We process two parallel diffusion branches: a positive branch that incorporates the desired shadow direction, and a negative branch that suppresses an undesired shadow direction (e.g., via an opposing light direction). Fig. 2 shows our dual-branch guidance mechanism. Classifier-free guidance then combines these branches:
where sets the guidance scale and is the object mask downsampled to the latent resolution. This dual-branch strategy reinforces the intended lighting on the object while naturally blending it with the background.
4.4 Final image synthesis
After the guided diffusion process, the resulting latent is decoded with the VAE decoder. The background preservation strategy of [78] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
5 Evaluation on 3D object compositing
Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene. To address this, we build over the evaluation setup of [78] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas. We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render. In this section, we describe how these scenarios are used for quantitative evaluation and a user study to assess the realism and controllability of the methods.
5.1 Evaluation dataset
We create datasets for the two scenarios using the Blender scene information graciously provided by the authors of [78]. We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [70]. We also improve the ground truth intrinsic maps by providing correct antialiased versions.
“Reference-based” scenario To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [78] to ensure lighting fidelity around the inserted object. We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [19]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3, in the “ground truth” column. This dataset is intended for quantitative evaluation (sec. 5.4).
“User-controlled” scenario This dataset simulates a scenario where a user specifies a desired light position. 8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images. This dataset is utilized in the user study (sec. 5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.
5.2 SpotLight specifics
As mentioned in sec. 3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [78] or RGBX [75]. Here, we provide details on ZeroComp, see supp. for RGBX.
We use the checkpoint trained on the Openrooms dataset [37] provided by the authors. The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions. Given its superior results, we adopt this method as our default model for 3D object compositing. To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [78], namely StableNormal [71] for normals, IID [32] for albedo, and ZoeDepth [8] for depth.
Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
Light-conditioned methods | |||||
DiLightNet [74] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
Neural Gaffer [29] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
Shadow-conditioned methods | |||||
IC-Light [76] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
ZeroComp+SDEdit [44] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
SpotLight | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
5.3 Baselines
We compare SpotLight with two types of baselines, that is methods that are conditioned on: the guiding shadow (like SpotLight); or an explicit lighting representation, through a environment map. We now describe per-method details. OpenImageDenoise [1] is also used to denoise the results of all methods to limit the impact of noise generated by diffusion models.
5.3.1 Shadow-conditioned methods
These methods, like ours, use the user-specified guiding shadow as input (obtained with EEVEE, see sec. 5.1).
ZeroComp [78]+SDEdit [44] We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map. To blend the shadows, we perform a refinement step, inspired by SDEdit [44], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
IC-Light [76] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning. We provide a rendering of the object lit by constant ambient lighting as input. No prompts are used.
5.3.2 Light-conditioned methods
Object relighting methods accept an explicit lighting representation as input in the form of a environment map, which we generate using the parametric light direction (see supp.). Unfortunately, these methods focus solely on object relighting—they do not create any cast shadows outside the object. To prevent this absence of shadows from severely hampering the realism of their results, we composite the guiding shadow onto the background image.
DiLightNet [74] DiLightNet requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [35] with the prompt “Putting aside the black background, this object is a”. DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [74]. The rendered object is then composited with the background.
Neural Gaffer [29] The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background. We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
Guiding shadow | IC-Light [76] | DiLightNet [74] | Neural Gaffer [29] | ZeroComp+SDEdit | Ours | Ground truth |
5.4 Experimental results
Tab. 1 presents quantitative results comparing the various baselines (c.f. sec. 5.3) with SpotLight (c.f. sec. 5.2) on our “reference-based” dataset, where the ground truth is available (sec. 5.1). Here, all methods are conditioned on the ground truth dominant light direction using their specific light parametrization (sec. 5.3). Fig. 3 shows qualitative results for each method. We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
(a) Overall realism study
(b) Lighting control study
5.5 User studies
Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines. In both cases, the Thurstone Case V Law of Comparative Judgement [61] is used to obtain a scaling -score for each method, where a higher scale indicates a higher human preference. Here, we give a high-level overview of those user studies. Refer to the supplementary material for additionnal details and studies.
User study I: Overall realism
Here, we use the “user-controlled” version of the dataset (see sec. 5.1) to evaluate the realism of results obtained by conditioning on a specific lighting direction, even if it is not aligned with the scene lighting conditions. We compare SpotLight against the three baselines with the best PSNR from tab. 1, namely Neural Gaffer [29], IC-Light [76], and ZeroComp [78]+SDEdit [44]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, as well as the left-right order. Note that we also add three sentinel images to discard observers that did not understand the task. Examples are shown in fig. 3.
User study I was completed by observers. The results, shown in fig. 4a, show that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
User study II: lighting control Since the goal of our method is to give full artistic control over the object relighting, we design a second two-alternative forced choice user study to evaluate lighting controllability separately from the realism. The users were shown two videos side by side and were asked to “Select the video where the light direction has the most impact on the lighting of the object, by considering both the appearance of the object and its shadows.”. The videos show a virtual object relit by a rotating light source (see supp.). Here, we compared the two methods with the highest realism according to user study I: SpotLight and ZeroComp+SDEdit. Since this task aims to evaluate the local relighting within the object mask only, we replace the shadowed background with the output from our method.
User study II was completed by observers. As seen in fig. 4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy.
5.6 Effect of parameter selection
Our method employs parameters that control the degree of shadow guidance ( in eq. 3) and the shadow blending ( in eq. 2). Tab. 2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the ground truth is available (c.f. sec. 5.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible (refer to the supp. for a qualitative comparison).
Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
, (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
(no guidance) | 31.69 | 0.976 | 0.029 | 0.011 | 0.029 |
28.68 | 0.966 | 0.043 | 0.015 | 0.036 | |
30.81 | 0.974 | 0.032 | 0.012 | 0.029 | |
29.24 | 0.969 | 0.039 | 0.014 | 0.034 |
No guidance () | (ours) |
6 Applications
This section highlights additional capabilities derived from our framework to further demonstrate its versatility.
Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and, in the supp., pixel height in 2D [58]. However, one can simply draw the desired shadow! Fig. 6 shows that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
Full image relighting As shown in fig. 7, we can extend SpotLight to achieve full-scene relighting. Although our method is not explicitly designed for this task, we find our shadow guidance to be effective in plausibly relighting the background. First, we relight the object using SpotLight with the desired guiding shadow. Next, we apply a version of ZeroComp trained with an inverted shading mask, using only the object shading to relight the background. We repeat this process for different light directions and average the resulting predictions to obtain an average relit background. Finally, we compute a per-pixel highlight map by dividing each background prediction by the average and multiplying it by the original background. This process ensures that background shadows and highlights consistently align with the object’s shadow.
7 Discussion
Incorrect user-defined shadow | Correct shadow |
Limitation: physically incorrect shadows Since SpotLight imposes no constraints on the input shadow, the user is free to specify a shadow direction that potentially contradicts the real lighting direction in the background image. In this case, SpotLight will still attempt to generate an image that looks as realistic as possible, given the shadow conditioning, as shown in fig. 8.
We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without requiring any additional training. Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
While the model can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
Acknowledgements
This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all members of the lab for discussions and proofreading help.