### v3

# SpotLight: Shadow-Guided Object Relighting via Diffusion

Abstract
Recent work has shown that diffusion models can serve as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise and controllable lighting can be achieved without any additional training, simply by supplying a coarse shadow hint for the object. Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, is entirely training-free and leverages existing neural rendering approaches to achieve controllable relighting. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting. We also demonstrate other applications, such as hand-scribbling shadows and full-image relighting, demonstrating its versatility.
| Background image | Previous method | SpotLight enables shadow-guided lighting control | |||
1 Introduction
Object relighting—the process of inserting a virtual object in an image and lighting it appropriately—has traditionally been confined to 3D graphics pipelines, where lighting is explicitly modeled and rendered. In contrast, object relighting in real photographs remains a fundamentally ill-posed problem since it requires inferring the physical properties of the scene: its geometry, materials, and illumination are not observed, and their complex interactions—such as specular reflections, soft shadows, or material-dependent reflectance—must be inferred or approximated.
Recent advances in image generative models, especially diffusion-based methods [rombach2022high], have opened up new possibilities for object relighting [zhang2025iclight, zhang2025zerocomp, liang2025diffusionrenderer]. These models encode strong priors over the appearance of scenes under varied lighting, allowing implicit reasoning about illumination, material, and geometry. Yet, precise and localized relighting control remains limited, particularly in the context of virtual object insertion, where a single object must be relit.
We focus on the problem of local lighting control for object compositing using generative models. Specifically, we aim to relight an inserted object such that the object conforms to a specified lighting direction and intensity around it, while blending it naturally with the background. Achieving such control is challenging: the relighting must align with physical cues like shadows and shading, yet remain flexible enough to work across varied materials, perspectives, and image contexts. Diffusion models offer powerful priors, but they are not inherently controllable without carefully designed guidance.
We present SpotLight, a training-free method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [zeng2024dilightnet, jin2024neural_gaffer]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer. Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it. We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object to insert in the image.
SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object. Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control. This mechanism is applied to recent diffusion models pre-trained to render images from input intrinsic maps (depth, normals, albedo, etc.) [zhang2025zerocomp, zeng2024rgb].
SpotLight: Shadow-Guided Object Relighting via Diffusion shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [eisemann2011realtime] according to the desired lighting direction. From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting.
In summary, our contributions are as follows. First, we introduce a framework that enhances pre-trained diffusion renderers with object relighting control without requiring additional training. We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models. Finally, we release an evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
2 Related work
Conditional image generation. Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [sohl2015deep, ho2020denoising, song2020denoising, rombach2022high]. These models offer control mechanisms for various modalities, such as text [ramesh2022hierarchical, rombach2022high], image content [saharia2022palette, gal2022image], intrinsic images [zhang2025zerocomp, luo2024intrinsicdiffusion, zeng2024rgb, liang2025diffusionrenderer], classifier guidance [dhariwal2021diffusion] and cross-modal data [brooks2023instructpix2pix, ye2023ip]. Notably, training-free guidance methods [meng2022sdedit, epstein2023diffusion, avrahami2022blended, avrahami2023blendedlatent, ho2021classifierfree] have enabled image editing capabilities without specific training, forming the basis of our approach.
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for harmonization [reinhard2001color, pitie2005n, lalonde2007using, tsai2017deep] do not prioritize physically accurate shading of the objects. Recently, [poirier2024diffusion] utilized a single-view multi-illumination dataset [murmann2019multi] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [xing2024retinex, bhattad2024stylitgan] for relighting indoor scenes. IC-Light [zhang2025iclight] enforces consistency in appearance to relight portraits and various objects. LumiNet [xing2025luminet] and ScribbleLight [choi2025scribblelight] achieve image relighting through illumination transfer or user-defined scribbles, respectively. Some methods specialize in outdoor scenes, using geometric priors like depth [griffiths2022outcast] or normalized coordinates [kocsis2024lightit] or both [lin2025urbanir]. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [bhattad2022cut] to harmonization of foreground-background albedo [careaga2023intrinsic]. Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [zhang2025zerocomp, zeng2024rgb, liang2025diffusionrenderer], which can be leveraged to achieve zero-shot object composition. However, these methods have limited lighting control. ZeroComp [zhang2025zerocomp] provides no lighting control, RGBX [zeng2024rgb] can only provide lighting control through a text prompt, and DiffusionRenderer [liang2025diffusionrenderer] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set. Other diffusion-based approaches, such as Neural Gaffer [jin2024neural_gaffer], DiLightNet [zeng2024dilightnet] and IllumiNeRF [zhao2024illuminerf] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. LightLab [magar2025lightlab] achieves fine-grained control over existing light sources, but does not handle object insertion. Careaga et al. [careagaRelighting] propose a sophisticated inverse rendering method for data generation, wheras SpotLight is training-free. IntrinsicEdit [lyu2025intrinsicedit] proposes a training-free method for lighting-based editing, but, unlike our approach, lacks explicit control over object-specific virtual lighting effects. We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
Explicit lighting estimation approaches infer HDR lighting from a single image [gardner2017sigasia, garon2019cvpr, li2020cvpr, zhu2022irisformer, phongthawee2024diffusionlight, liang2024photorealistic], though they generally lack controllability. Parametric models [gardner2019iccv, weber2022editable, dastjerdi2023everlight] or GAN inversion methods [wang2022stylelight] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
Shadow generation. Generative models for rendering realistic shadows [zhang2019shadowgan, liu2020arshadowgan, hong2022shadow, liu2024shadow], or harmonizing rendered shadows with the image [valenca2023shadow] have been proposed. Other methods enable controllable shadows for 3D [sheng2021ssn] and 2D objects [sheng2022controllable, sheng2023pixht]. Our approach can leverage existing shadow generation methods (e.g., [tasar2024controllable]) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
Image intrinsics. Decomposing images into albedo and shading has long been studied [barrow1978recovering, kovacs2017shading, philip2019multi, wu2023measured, careaga2023intrinsic]. Numerous works have also focused on estimating scene geometry, for example, depth [ranftl2020towards, bhat2023zoedepth, yin2023metric3d, ke2024repurposing, yang2024depth, depth_anything_v2] and normals [eftekhar2021omnidata, bae2021estimating, ranftl2021vision, bae2024rethinking, ye2024stablenormal]. Recently, approaches such as [li2020cvpr, zhu2022irisformer] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [bhattad2024stylegan, du2023generative, ke2024repurposing, ye2024stablenormal, anonymous2024intrinsiccontrolnet, luo2024intrinsicdiffusion]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
| Guiding shadow | IC-Light [zhang2025iclight] | DiLightNet [zeng2024dilightnet] | Neural Gaffer [jin2024neural_gaffer] | ZeroComp+SDEdit | Ours | Ground truth |
3 Background: diffusion renderers
Recent work has shown that diffusion models can be adapted to generate photorealistic images from intrinsic maps including materials (e.g., albedo, roughness, metalness), geometry (e.g., surface normals and depth), and shading. This capability was built using a ControlNet in ZeroComp [zhang2025zerocomp] or full finetuning in RGBX [zeng2024rgb]. We refer to this new class of conditional generation models as “diffusion renderers”.
Specifically, ZeroComp [zhang2025zerocomp], which we will use in sec. 5, is trained to generate images from albedo, normals, depth, and partially-masked shading. It can be used for compositing in a zero-shot manner by alpha-compositing the intrinsics of the background and the object, and by masking out a region on and around the object in the shading map. The network then generates the full image with realistic shading. The intrinsics of the object are rendered using simple graphics shaders, while those of the background are estimated using pre-trained networks (e.g., [bhat2023zoedepth] for depth, see sec. 5.2). In the case of RGBX [zeng2024rgb], their XRGB network achieves object insertion by accepting as input alpha-composited albedo, normals, and a masked image. They then train an inpainting version of their network to generate the final image. Unlike traditional rendering, these methods offer no control over the lighting conditions.
In this work, we use such a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, without any additional training. We now describe our approach in detail.
4 SpotLight
We employ a user-provided, guiding shadow to steer a pre-trained diffusion renderer (c.f. sec. 3) to follow the desired lighting direction. Our key insight is that even an approximate shadow encapsulates enough important lighting information and, by integrating this cue in the denoising process, we can simultaneously refine the composite appearance and its shadow, yielding a natural blend with the scene.
4.1 Approach overview
An overview of SpotLight is illustrated in fig. 2. In addition to the intrinsic maps for the backbone diffusion renderer, our approach requires two inputs: an object mask and a user-specified guiding shadow which guides the latent diffusion process. The guiding shadow can be obtained in several ways—in this paper, we experiment with fast rasterization [eisemann2011realtime], a trained shadow generation model [tasar2024controllable], and hand-drawn scribbles. Without retraining the diffusion renderer, we steer its generative process by incorporating the guiding shadow into the latent space.
4.2 Blending shadows with the background
Provided the guiding shadow , we first need to ensure that the diffusion model will properly blend the shadow with the background. We take inspiration from Blended Latent Diffusion [avrahami2023blendedlatent] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadow roughly composited over the image, . In practice, we obtain by compositing the object albedo and the guiding shadow on the background. The update made at each timestep to incorporate the desired shadow latents is
where is the shadow latent weight and is the shadow mask, bilinearly downsampled to the latent resolution (), with the edge opacity increased and dilated to maintain the desired shadow softness (see supp. for details). Noise is added to the encoded composite using the DDIM [song2020denoising] noise scheduler. Different from [avrahami2023blendedlatent], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
4.3 Enhancing lighting control
Simply running a diffusion renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec. 5.6). We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by . Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample. From the outputs of the diffusion renderer (typically, -predictions [salimans2022progressive], i.e., linear combination of noise and estimated image), we then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After eq. 2, we run a regular DDIM sampling step to obtain the latents .
4.4 Final image synthesis
After the guided diffusion process, the resulting latents are decoded with the VAE decoder. The background preservation strategy of [zhang2025zerocomp] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
5 Evaluation on 3D object compositing
Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene. To address this, we build over the evaluation setup of [zhang2025zerocomp] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas. We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render. In this section, we describe how these scenarios are used for quantitative evaluation and a user study to assess the realism and controllability of the methods.
5.1 Evaluation dataset
We create datasets for the two scenarios using the Blender scenes graciously provided by the authors of [zhang2025zerocomp]. We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [depth_anything_v2]. We also improve the ground truth intrinsic maps with antialiased versions.
“Reference-based” scenario. To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [zhang2025zerocomp] to ensure lighting fidelity around the inserted object. We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [eisemann2011realtime]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3, in the “ground truth” column. This dataset is intended for quantitative evaluation (sec. 5.4).
“User-controlled” scenario. This dataset simulates a scenario where a user specifies a desired light position. 8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images. This dataset is used in the user study (sec. 5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.
5.2 SpotLight specifics
As mentioned in sec. 3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [zhang2025zerocomp] or RGBX [zeng2024rgb]. Here, we provide details on ZeroComp, see supp. for RGBX.
We use the variant trained on Openrooms [li2020openrooms] provided by the authors. The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions. Given its superior results, we adopt this method as our default model for 3D object compositing. To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [zhang2025zerocomp], namely StableNormal [ye2024stablenormal] for normals, IID [kocsis2023iid] for albedo, and ZoeDepth [bhat2023zoedepth] for depth.
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| Light-conditioned methods | |||||
| DiLightNet [zeng2024dilightnet] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
| Neural Gaffer [jin2024neural_gaffer] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
| Shadow-conditioned methods | |||||
| IC-Light [zhang2025iclight] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
| ZeroComp+SDEdit [meng2022sdedit] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
| SpotLight | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
5.3 Baselines
We compare SpotLight with two types of baselines, that is methods that are conditioned on: the guiding shadow (like SpotLight); or an explicit lighting representation, through a environment map. We now describe per-method details. OpenImageDenoise [OpenImageDenoise] is also used to denoise the results of all methods to limit the impact of rendering noise present in the synthetic training data of some models.
5.3.1 Shadow-conditioned methods
These methods, like ours, use the user-specified guiding shadow as input (obtained with EEVEE, see sec. 5.1).
ZeroComp [zhang2025zerocomp]+SDEdit [meng2022sdedit]. We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map. To blend the shadows, we perform a refinement step, inspired by SDEdit [meng2022sdedit], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
IC-Light [zhang2025iclight]. We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning, with the object lit by a constant ambient lighting as input. No prompts are used.
5.3.2 Light-conditioned methods
Object relighting methods accept an explicit lighting representation as input in the form of a environment map, which we generate using the parametric light direction (see supp.). Unfortunately, these methods focus solely on object relighting—they do not create any cast shadows outside the object. To prevent this absence of shadows from severely hampering the realism of their results, we composite the guiding shadow onto the background image.
DiLightNet [zeng2024dilightnet] requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [li2023blip2] with the prompt “Putting aside the black background, this object is a”. DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [zeng2024dilightnet]. The rendered object is then composited with the background.
Neural Gaffer [jin2024neural_gaffer]. The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background. We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
5.4 Experimental results
Tab. 1 presents quantitative results comparing the various baselines (c.f. sec. 5.3) with SpotLight (c.f. sec. 5.2) on our “reference-based” dataset, where the ground truth is available (sec. 5.1). Here, all methods are conditioned on the ground truth dominant light direction using their specific light parametrization (sec. 5.3). Fig. 3 shows qualitative results for each method. We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
(a) Overall realism study
(b) Lighting control study
5.5 User studies
Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines. In both cases, the Thurstone Case V Law of Comparative Judgement [thurstone2017] is used to obtain a -score for each method, where a higher value indicates higher human preference. Here, we give an overview of those user studies. Refer to the supp. for additional details and studies.
User study I: Overall realism
Here, we use the “user-controlled” version of the dataset (see sec. 5.1) to evaluate the realism of results obtained by conditioning on a specific lighting direction, even if it is not aligned with the scene lighting conditions. We compare SpotLight against the three baselines with the best PSNR from tab. 1, namely Neural Gaffer [jin2024neural_gaffer], IC-Light [zhang2025iclight], and ZeroComp [zhang2025zerocomp]+SDEdit [meng2022sdedit]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, and the left-right order. Three sentinel images were added to discard observers that did not understand the task. Examples are shown in fig. 3.
User study I was completed by observers. Fig. 4a shows that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
User study II: lighting control Since the goal of our method is to give full artistic control over the object relighting, we design a second two-alternative forced choice user study to evaluate lighting controllability separately from the realism. The users were shown two videos side by side and were asked to “Select the video where the light direction has the most impact on the lighting of the object, by considering both the appearance of the object and its shadows.”. The videos show a virtual object relit by a rotating light source (see supp.). Here, we compared the two methods with the highest realism according to user study I: SpotLight and ZeroComp+SDEdit. Since this task aims to evaluate the local relighting within the object mask only, we replace the shadowed background with the output from our method.
User study II was completed by observers. As seen in fig. 4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy. Please see the supp. for more user study results and experimental details.
5.6 Effect of parameter selection
Our method uses two tunable parameters: the local guidance scale ( in eq. 2) and the shadow blending weight ( in eq. 1). Tab. 2 reports metrics obtained on the “reference-based” dataset where the ground truth is available (c.f. sec. 5.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible. Please refer to the supp. for a qualitative comparison for the , and the shadow softness (light radius) parameters.
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| , (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
| (no guidance) | 31.69 | 0.976 | 0.029 | 0.011 | 0.029 |
| 28.68 | 0.966 | 0.043 | 0.015 | 0.036 | |
| 30.81 | 0.974 | 0.032 | 0.012 | 0.029 | |
| 29.24 | 0.969 | 0.039 | 0.014 | 0.034 |
| No guidance () | (ours) |
6 Applications
This section highlights additional capabilities derived from our framework to further demonstrate its versatility.
Scribbles as shadows. Instead of relying on shadow mapping in 3D, one can simply draw the desired shadow! Fig. 6 shows that SpotLight can realistically relight the object and refine its shadow even when it is a user-drawn scribble.
2D objects. SpotLight can be applied to real 2D objects from existing photographs by estimating their intrinsics, as in sec. 5.2, and generating desired shadows with shadow generation methods such as [tasar2024controllable], as illustrated in fig. 7.
| Input object | Relit object on target background | ||
Reflective materials. When using a ZeroComp backbone pre-trained on the InteriorVerse dataset [zhu2022learning], SpotLight naturally generalizes to metallic objects and generates controllable specular reflections on the object, see fig. 8.
Additional relighting results. SpotLight also shows other capabilities: relighting for more diverse objects and outdoor scenes, simulating multiple light sources by combining outputs with different guiding shadows, and full-image relighting. Due to space constraints, the implementation details and results are provided in the supp.
7 Discussion
| Incorrect user-defined shadow | Correct shadow | ||
Limitation: physically incorrect shadows. Since SpotLight imposes no constraints on the input shadow, users may specify a shadow direction that potentially contradicts the real lighting in the background. In this case, SpotLight still attempts to render a visually plausible composition consistent with the given shadow, as shown in fig. 9.
We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without any additional training. Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
While our method can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
Acknowledgements. This research was supported by FRQNT scholarship 328810, NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all lab members for discussions and proofreading help.